How much energy is associated with a mole (6.022 x 1023) of photons with a frequency of 4.30 x 1014 s-1 ( Answer in kJ to 3 sig figs)
Answer #1

Energy of photons E= nh\nu - -------------- (1)

n is number of photons,

h is plank's constant, 6.626 * 10-34 J s

v is frequency of light.

According to question,

n = 6.02* 1023

v = 4.30 * 1014 s​​​​​​-1​

Substituting respective values in equation (1)

E = 6.02 * 1023 * 6.626 * 10-34 J s * 4.30 * 1014 s​​​​​s-1

E = 171577.62 J = 171.578 kJ. (I kJ = 1000J)